Fermi National Laboratory


Accelerator Update

Friday July 9
The day shift began with the Tevatron (TeV) in shot setup, and with MiniBooNE taking beam.

At 11:30 AM, a Booster RF station (BRF16) tripped for the first of six times in twenty minutes. Operations bypassed the station while experts investigated. They had the station back online twenty minutes later.

Operations established store 3639 at 11:51 AM with an initial luminosity of 76.85E30.

Antiproton Source (Pbar) experts conducted about four minutes of flying wire studies in the Accumulator and then resumed stacking at 11:55 AM.

Operations established beam to the Meson MTest experiment at 12:24 PM. Then at 12:31 PM, the experimenters decided they didn't want any more beam. Operators ran the beam to the Meson dump.

Operations began stashing antiprotons into the Recycler at 1:02 PM. The process took about fifty-three minutes.

A Linac RF station (KRF1) tripped off at 1:19 PM due to a flow switch. Linac technicians applied some mechanical persuasions to the flow switch and the switch understood that it was in its own best interest to begin working again.

Pbar resumed stacking at 1:55 PM.

Operations halted stacking and prepared the Pbar tunnels for access at 2:05 PM.

At 2:25 PM, a Booster power supply (EAPS) tripped off along with an RF station (BRF2). Later it was discovered to be due a timeline (TLG) problem and was fixed.

At 2:32 PM, A TeV RF station (TRF2) tripped off due to a cavity water heater problem. Operators notified experts.

Operations reestablished beam to the MTest experiment at 3 PM.

At 4:18 PM, a tech discovered a pinhole LCW leak on a Pbar Debuncher magnet. This required brazing to repair; experts approved this work.

At 4:43 PM, Operations reported that experts were working on understanding all the MIRF trips from Thursday.

Around 5 PM, an MI expert began a Mixed Mode study for sending 120 GeV beam to Switchyard. The study ended at 6:02 PM.

The water techs reported at 6:28 PM that the brazing on the Pbar Debuncher magnet was completed.

A NuMI study, which had started this afternoon, ended at 6:35 PM.

At 6:37 PM, Operations reported that a Booster Debuncher dipole bus (D:IB) tripped off while cycling. Operators investigated and called an expert for advice. D:IB was back on by 7:07 PM. The expert thought it was a sequencer problem.

Water techs reported that they had finished installing a new LCW return hose on an Accumulator magnet (A2B3) at 7:09 PM. There was also a shunt spraying water at D106 repaired.

Pbar resumed stacking at 7:28 PM.

MTest experimenters said that they were finished using beam at 9:40 PM.

At 10:15 PM, the MI abort kickers fired. The system reset okay.

Saturday July 10
The midnight shift began with Operations monitoring store 3639 and the Recycler stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations established beam to the Meson MCenter experiment at 3:06 AM.

Operations established beam to MTest experiment at 7:10 AM.

At 8:48 AM, MCenter said that they were through with the beam.

The Pbar lithium lens (D:LN) tripped off on two indications of a capacitor overvoltage and a ground fault at 9:08 AM. It reset okay. It tripped off again twenty minutes later. Operators reset the lens and called an expert.

At 9:50 AM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum pump tripping off. Everything reset okay.

At 11:01 AM, a Booster RF station (BRF1) tripped off. It reset okay.

At 1:27 PM, a Booster RF station (BRF14) tripped off. It reset okay.

At 4:44 PM, BRF1 tripped off. It reset okay.

Operations began conducting end-of-store studies at 6:56 PM.

At 7:23 PM, BRF1 tripped off. It reset okay.

Operations terminated store 3639 at 7:32 PM.

At 8:09 PM, Operations reported that a MI RF station (MIRF13) had tripped off a few times on gap envelope faults. Operators investigated and then at 9:05 PM called in an expert.

The RF expert arrived at 10 PM and had MIRF13 operating normally by 11:25 PM.

At 10:33 PM, a TeV RF station (TRF8) tripped off on repeated envelope gaps. The RF expert had the station back up by 11:50 PM.

Sunday July 11
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Operations established store 3641 at 12:58 AM with an initial luminosity of 79.3E30. Pbar resumed stacking.

At 2:06 AM, a Pbar power supply (PMAG) tripped off on a ground fault indication. Operators investigated and at 2:52 AM reported that the supply would trip on a ground fault after a few pulses. Operations called in an expert.

The expert investigating the PMAG supply arrived around 3:30 AM and reported at 5:20 AM that the supply was okay, but that there was a ground fault on the magnet. Operations prepared Pbar for an access.

Operations established beam to MTest at 6:50 AM.

At 7:24 AM, Operations reported that MIRF13 had tripped off three times. It reset okay each time.

At 7:56 AM, Operations reported that MIRF13 was off and bypassed. The MI RF system was reconfigured to work without this station.

The Booster EAPS tripped at 8:21 AM. It reset okay.

At 9:46 AM, BRF7 & 8 tripped off due to a vacuum pump tripping off. Everything reset okay.

At 10 AM, Operations reported that a SY SEPTA (S:FSEPV) had been tripping off and then turning itself back on.

The Booster EAPS tripped at 10:28 AM. It reset okay.

At 10:30 AM, Operations reported a ComEd power glitch. Three TeV RF stations tripped off and that caused store 3641 to abort. ComEd reported that a transformer in Lombard had blown-up. The good news was that the stack and stash survived.

Operations reestablished beam to MiniBooNE at 10:34 AM.

A TeV expert began a study period at 12:09 PM.

Pbar experts reported at 12:28 PM that techs were beginning to install the new PMAG.

MTest experimenters said at 12:44 PM that they were done with beam for a while.

Pbar experts reported that the PMAG test pulsing began at 2:33 PM.

Operations established beam to the Pbar Accumulator at 2:50 PM.

At 2:53 PM, a Linac RF station (KRF5) tripped off. It reset okay.

Operations loaded a 12x0 study-store for TeV experts at 3:05 PM.

At 3:26 PM, KRF5 tripped off again. It took a little while, but eventually it reset okay.

At 3:32 PM, a TeV RF station (TRF8) tripped off due to a gap envelope problem.

Operations reestablished beam to MTest at 3:39 PM.

Operations loaded another 12x0 study-store for TeV experts at 4:18 PM, at 5:39 PM, and then again at 7:34 PM.

At 5:05 PM, MIRF7 tripped off and kept tripping off. An expert investigated the problem.

Monday July 12
The midnight shift began with TeV experts conducting studies, with Pbar stacking, with Operations monitoring the stash, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 2:05 AM and at 4:03 AM established store 3650 with an initial luminosity of 66.3E30.

Operations established beam to MCenter at around 5:30 AM.

Operations set up Recycler to stash antiprotons at 7:18 AM. The first transfer didn't make it because two power supplies (HV200 & HV202) tripped off on phase unbalance. These two supplies tripped off several more times and caused Operations to cancel adding any antiprotons to the existing stash. Pbar resumed stacking.

Plans for Monday Morning
The plans for today are to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/12/2004   email Fermilab